# Jyler Menard,a and RA Mansbacha

### v2

@@ 2.2 Latent Bayesian Optimization
Latent Bayesian Optimization attempts to resolve the issue of optimizing over large, discrete design spaces, such as, in our case, amino acid sequences describing peptide candidates. Rather than attempting to search for optimal peptide sequences in the discrete combinatorial space, we instead perform the optimization over a continuous-valued latent space, where we have access to both an encoder $E:X\rightarrow Z$ encoding sequences into a latent space, and a decoder $D:Z\rightarrow X$ that reconstructs the sequence corresponding to a latent point. Here $Z$ is typically the latent space of a generative model. The conventional choice is to use the latent space of a VAE, where the encoded mean $\mu_{\theta}(x)$ of a trained VAE with fitted model weights $\theta$ is used in the optimization 22. More formally, Latent Bayesian Optimization is the process of solving
$\mathop{\text{argmax}}_{x\in X}\vec{f}(x)$
@@
$\mathop{\text{argmax}}_{z\in Z}\vec{f}(D(z))$
i.e. Bayesian Optimization over the latent space of a generative model or other invertible mapping. The surrogate model $\hat{f}:Z\rightarrow\mathbb{R}$ approximates the relationship between latent points and their corresponding objective value; for a peptide sequence $\vec{x}$ and its encoded representation $\vec{\mu}_{\theta}(\vec{x})$, the surrogate model approximates $\vec{\mu}_{\theta}(\vec{x})\rightarrow\vec{f}(\vec{x})$. At iteration $l$, the acquisition function determines the subsequent point in the latent space $\vec{\mu}_{l+1}$ to decode into a peptide sequence $\vec{x}_{l+1}$ we can evaluate with the objective function $\vec{f}(\vec{x}_{l+1})$. In other words, rather than performing the Bayesian Optimization on what we want to design (peptide sequences) we perform it over a (learned) representation of those sequences.
@@
In this work we are interested in finding peptide sequences with strong antimicrobial activity. We choose to use minimum inhibitory concentration (MIC) as a measure for antimicrobial activity. The MIC refers to the smallest concentration of peptide that inhibits the growth of bacteria in a standardized assay31. We make this choice because it is a theoretically continuous, standard, and meaningful measure of antimicrobial activity, and standardized data is available for measurements on it 32. Because a lower MIC corresponds to stronger antimicrobial activity, we maximize the negative of MIC: $M=-\log_{10}(MIC)$, thereby minimizing MIC.
In this pilot study, we wished to assess the impact of algorithmic choices and ensure our understanding of the design space prior to engaging in time-consuming experimental work. Therefore, we use a proxy model (also called an “oracle”) based on traditional machine learning techniques (see next section) as our “ground truth” to estimate the MIC values of sequences for which no experimentally-measured MIC is available. This is analogous to proof-of-concept small-molecule discovery projects that optimize for QED or calculated LogP; a discipline-relevant proxy model is used to motivate an algorithmic exploration. In future work, we plan to leverage the procedure here to inform peptide design in more realistic contexts.
2.3.1 Support Vector Regression Oracle
To serve as an oracle to rapidly evaluate our “true” objective function, we trained a Support Vector Regression (SVR) model on data compiled by Witten & Witten (2019)32 by using the scikit-learn implementation 33. For the outputs of the oracle, we employed the base-10 logarithm of the MIC values $\log_{10}(\text{MIC})$ to ensure the values are distributed more symmetrically (Fig. S2). For the inputs to the oracle, we used easily-computed physico-chemical descriptors. We used a previously-validated feature selection procedure34 to identify the most relevant subset from a larger set of possible descriptors.
Computing all descriptors available in the ProPy3 Python package 35 resulted in 1529 features per sequence. We reduce the number of features necessary for the model via a two-step feature selection procedure. We perform an initial feature screening stage using the minimum Redundancy Maximum Relevance (mRMR) algorithm36, in which one attempts to identify those features that have the least overlap that are the most relevant to the prediction of the output of interest. We defined redundancy as the mutual information between pairs of physico-chemical features and relevancy as the mutual information between features and predicted outputs (i.e. $\log_{10}(\text{MIC})$).
[…]
@@ 2.4 Variational Autoencoder
For the generative model whose associated latent space becomes our continuous design space of peptide sequences, we employ a variational autoencoder (VAE) based neural network. A VAE is a probabilistic deep learning architecture comprising two neural networks: the encoder $E_{\phi}:X\rightarrow Z$, which takes an input sequence and projects it to a continuous-valued latent space $z=E_{\phi}(x)$; and the decoder $D_{\theta}:Z\rightarrow X$ modules, which maps a point in the latent space back to the original input sequence space $\hat{x}=D_{\theta}(z)$ with $\phi$ and $\theta$ representing learned model weights 39. Once the model has been adequately trained, we can use the latent space and the decoder to generate new peptides, sampling from a distribution that is similar to the distribution found in the training set.
Training a VAE entails minimizing the Evidence Lower Bound (ELBO), a combination of a reconstruction loss, which measures the difference between the input sequence and the corresponding generated output sequence, and the Kullback-Leibler Divergence which functions to push the encoder probability distribution to be close to a Gaussian prior modeled by the latent space. We define the reconstruction loss as is typically done for sequence-based tasks 40 as the cross-entropy of the probability of predicting the next token $x_{i+1}$ conditioned on a context $\vec{c}$, usually the previous tokens in the sequence:
@@
We use a VAE with a transformer architecture as described previously 41, 21, known therein as a TransVAE. The transformer architecture is largely similar to the architecture originally proposed in Vaswani et al. 42 It starts with an embedding layer and a positional encoding layer then together feed into the encoder module consisting of self-attention blocks followed by convolution blocks. Each self-attention block is a sequence of layers: layer normalization, multi-head self-attention, dropout, a residual connection, then layer normalization, feed-forward, dropout, and another residual connection. After the self-attention blocks, there is a convolution block consisting of three layers of convolution, pooling, and leaky ReLU. The convolution block outputs the mean and variance for the VAE latent space. The decoder module consists of largely the same architecture in reverse: convolution blocks followed by masked self-attention blocks. In our previous results 21 we found evidence suggesting the TransVAE architecture may limit latent space interpretability compared to other architectures, however we did observe high performance at the assigned tasks. Here, we chose to use a transformer architecture because of their effectiveness at sequence-to-sequence tasks, and because traversing their latent space can correlate well with a predicted property 21, an important aspect of this work. Additionally, to limit the scope of this work, we focused on only the TransVAE architecture, but in future work will expand to other architectures and representation learning choices.
As previous work has shown, jointly training a property predictor can induce latent space organization 22, 21. In this work, we investigate the impact of the choice of predictive task and the availability of data associated with that task on the performance of the LBO algorithm. To do so, we train multiple TransVAE models jointly with a property predictor consisting of two feed-forward layers (see next section). The property predictor loss term $\mathcal{L}_{\text{PropPred}}$ is a simple mean-squared error between the property predictor’s outputs $y_{\text{pred}}$, and the computed property values $y_{\text{true}}$,
@@ 2.5 TransVAE Models and Training
To investigate the importance of additional information on the LBO algorithm and on the latent space that it traverses, we trained TransVAE models in conjunction with property predictors of different combinations of physico-chemical characteristics. Jointly training property predictors requires target values of these characteristics for corresponding (latent representations of) sequences. We computed Boman Index, net charge at a pH of 7.2, and hydrophobicity of every sequence in both the training set and test set. Boman Index is defined by summing the solubility values of individual amino acids in the peptide, normalized by length. It functions as an estimate of whether a peptide will bind to membranes or other proteins. For hydrophobicity, we use the Kyte-Doolittle scale and average the hydrophobicity values of amino acids in a given peptide for an overall estimate of peptide affinity with water. Charge at a pH of 7.2 is simply computed by summing the predicted charge of each amino acid in the peptide at a pH of 7.2. These physicochemical properties are of interest in AMP design because of they relate to whether an AMP will interact with or penetrate a cell membrane.
@@
Overall, we studied the following models: (i) TransVAE trained jointly with a regressor directly predicting oracle values, that is, predicting predicted-log10mic values; (ii) TransVAE trained jointly with a regressor predicting Boman index values; (iii) TransVAE trained jointly with a regressor predicting charge values; (iv) TransVAE trained jointly with a regressor predicting hydrophobicity values; (v) TransVAE trained jointly with a regressor predicting both Boman index and charge; and (vi), TransVAE trained jointly with a regressor predicting Boman Index, charge, and hydrophobicity values.
@@ 2.6 Dataset
To train our TransVAE models we used a peptide sequence dataset consisting of 665,772 sequences. We sourced sequences from from AMP-specific databases DRAMP V3.0 44, APD3 45, GRAMPA 32, dbAMP 2.0 46, StarPep 47, and general protein databases SwissProt and TREMBL 48. Using peptide sequences from databases of known antimicrobial peptides ensures that the data distribution — and therefore the distribution being learned by the TransVAE — contains examples of sequences that are AMPs. Additionally, to ensure the training distribution also contained non-AMPs, we excluded those peptides from SwissProt annotated with the keywords antifungal, antimicrobial, antibiotic, antiviral defense, antiviral protein, secreted; although the impact of keyword combinations on predictive model performance has been discussed elsewhere49. Lastly, we clustered TREMBL sequences at 90% identity using CD-HIT 50, 51, then randomly sampled 7500 peptides at each length, yielding a subset of 668,343 peptides (the smallest length in the TREMBL set after clustering was 15). Once we removed duplicates, the final dataset of unique peptides from every database was 665,772 sequences. This final amount was split into a training set of 523,848 sequences and a test set of 141,924 sequences where we ensured that the training set and test set had the same proportion of verified AMP and verified non-AMP peptides (SI Fig. S1).
@@ 2.8 Latent Bayesian Optimization Procedure
To perform latent BayesOpt, we use the BoTorch Python package 28, leveraging its SingleTaskGP Gaussian Process object with input unit hypercube and output standard normalization, as suggested by the codebase. As noted above, we use the Log Expected Improvement (LogEI) acquisition function, which offers improved numerical stability compared to conventional Expected Improvement 30. For each BayesOpt experiment, we perform five different optimization runs of 500 iterations (’oracle’ calls). Each run is initialized with 100 points randomly sampled from predicted-$\log_{10}$(MIC) values (see Supplementary Information for additional details). Every BayesOpt experiment uses the same 5x100 initialization points, unless otherwise specified.
@@ 3.1 Property prediction leads to latent space organization, even in data-sparse conditions
To assess the latent spaces associated with the different models, we quantified the extent to which the top five principal components of a PCA projection correlated with different physicochemical properties used in training. We then employed these PCs to visualize relevant two-dimensional slices of the latent space. To do this we first computed each property over the training set of sequences. Next, we embedded those sequences in the full latent space of each model. Then we decomposed the space into its top five principal components. For each principal component, we computed the Pearson correlation coefficient ($r$) between it and the property values. We observed that sequence length was most correlated with the first principal component across 21/27 of the models (SI Fig S4g), with the second or third principal component being correlated with the property used during training to organize the latent space (SI Fig. S4h for Boman index; SI Fig. S4i for charge; SI Fig. S4j for hydrophobicity; SI Fig. S4k for oracle). Although we may desire the first principal component to be most correlated with the organizing property in most cases, we note that the top five principal components have similar amounts of explained variance, and that they capture on the order of $20\%$ of the total variance (SI Fig. S3). Additionally, previous work found sequence length can be an easy-to-learn pattern21. Patterns that are straightforward to recognize may become unexpectedly strongly correlated with a PC direction.
In Fig. 1, we visualize the PC components most highly correlated with Boman Index, Charge, and Hydrophobicity for the models trained with all three. In line with previous results 22, 21, we observe latent space organization when jointly-training a decoder and a property predictor. Additionally, training a property predictor predicting multiple different properties results in an organized latent space where different properties correlate most with different pairs of dimensions (Fig. 1 rows; SI Table S3, particularly Boman index and charge). Furthermore, varying percentage of property labels — thereby limiting the amount of labels the model has access to in training — still results in latent space organization, as indicated by the persistence of an orderly color gradient between the first and last rows of Fig. 1. Overall, increasing the number of properties and masking even a large proportion of the information for the predictor did not prevent the latent space from being organized.
Given that for 64-dimensional latent spaces we are visualizing just two-dimensional slices through PCA projections, it is possible that the organization is being induced by the projection. To check the extent to which PCA is distorting the original, high-dimensional manifold, and thus the extent to which it may be providing an illusion of organization when the high-dimensional cloud of points is not actually organized, we compute four manifold distortion metrics: trustworthiness and continuity 53, steadiness and cohesiveness 54.
@@
We observe similar levels of distortion between high-dimensional latent spaces and their PCA projections for models trained with different organizing properties and different percentages of labels (Fig. 2). Trustworthiness, continuity, and steadiness are all relatively high, with each organizing method exhibiting values >0.75>0.75 (Fig 2abc), suggesting high-quality projections with few artificial clusters (trustworthiness), minimal loss of clusters (continuity), and little stretching of distances between points (steadiness). Cohesiveness, lying near $0.6$ for all, is moderate, suggesting that compression is occurring (Fig 2d), as it necessarily must when moving from high to low dimensions. The similarity of the results and the comparatively low-to-moderate distortions observed provide confidence that the visualizations reflect extant trends between, and patterns in, the high-dimensional latent spaces.
@@
Given the number of different physicochemical properties by which a latent space may be organized, we perform an ablation study testing different numbers and combinations. We hypothesized that organizing the latent space with more properties would lead to peptide representations more amenable to BayesOpt. We summarize the performance at different iterations in Figs S10 and Figs 4, 5, where we investigate average best scores found under different conditions. We primarily focus on PCA5 models. We choose the top five components to balance capturing a larger amount of explained variance (near 20%, SI Fig. S3) with losing interpretability, and because three of the four scenarios where we varied the number of PCA components demonstrated five components performing better than two components (previous section).
@@
We consider single property organization first. There is no clear advantage of any single property when searching through the 64-dimensional latent spaces–the best-performing single-property models compared to other models with only organizing property changed are id-c-100%, id-b-75%, id-h-50%, id-h-2% at 100 iterations (Fig. 4), and id-b-100% and id-c-100%, id-b-75%, id-d-50%, id-c-25%, and id-b/c-2% at 500 iterations (Fig. 5). For PCA-projected spaces, we observe that charge-organized spaces tend to demonstrate better performance: at 100 iterations, PCA5-c-100% clearly outperforms b/h-100%, PCA5-c-75% clearly outperforms b/h-75%, PCA5-c-50% clearly outperforms b/h-50%, and PCA5-c-2% clearly outperforms b/h-2%. Only for the PCA5-25% models is charge not the clear winner, and in that case, it performs nearly as well as PCA5-h-25%. At 500 iterations, PCA5-c-100% clearly outperforms b/h-100%, PCA5-c-75% clearly outperforms b/h-75%, and PCA5-c-50% clearly outperforms b/h-50%. However, PCA-5-c-25% performs worse than b/h-25%, and all PCA5-2% models perform roughly the same, displaying rather poor performance at <0.3 average best scores. Thus, charge is a better organizing property than Boman index or hydrophobicity, particularly when more labels are available. This trend matches with a similar trend in relevancy: charge correlates more strongly and has higher mutual information with the oracle values when looking at the full peptide training set (SI Section S4.2).
@@
Considering multi-property organization, we still do not see clear trends in when searching through the full 64-dimensional latent space. After 100 iterations (Fig. 4), id-bch-100% is one of the best performers (with performance about equivalent to id-c-100%), and id-bch-75% outperforms other models, but the highest-performing model for lower label percentages is one or more of the single-property models. After 500 iterations (Fig. 5), we see similar model performance across the board, though id-bch-100% is one of the highest-performing models (tied with id-b-100%) and id-bc-75% is one of the highest-performing models (tied with id-b-75%) but for lower label percentages single-property models are the best, and the performance is not too dissimilar within error bars.
@@
When searching through a PCA projection, there is not a clear advantage to using multi-property organization. After 100 iterations (Fig. 4), at 100% and 75% label percentage, charge clearly outperforms multi-property organization. At 50% label percentage, PCA5-bc-50% outperforms PCA5-c-50%, at 25% label percentage, PCA-bch-25% by far outperforms any other models, and at 2% the highest performing model is PCA5-bch-2%. After 500 iterations (Fig. 5), at 100% and 75% label percentage, charge still outperforms multi-property organization. For lower label percentage, multi-property organization does outperform single-property organization (PCA5-bc-50% > PCA5-bch-50% > PCA5-c-50%, PCA5-bch-25% is the clear best performer, PCA5-bch-2% outperforms PCA5-bc-2% and both clearly outperform the single-property organizers), but neither bc nor bch is consistently better than the other. This suggests that multi-property organization may be better at lower label percentage, an observation which is supported by the fact that the best-performing model we observed in the study was PCA-20-oracle-2%, with a final performance of $\langle\mathcal{M_{\textrm{final}}}\rangle=$0.896 $\pm$ 0.092, where the oracle is itself a multi-property organizer, although it was not part of the ablation study.

### v3

@@ 2.2 Latent Bayesian Optimization
Latent Bayesian Optimization attempts to resolve the issue of optimizing over large, discrete design spaces, such as, in our case, amino acid sequences describing peptide candidates. Rather than attempting to search for optimal peptide sequences in the discrete combinatorial space, we instead perform the optimization over a continuous-valued latent space, where we have access to both an encoder $E:X\rightarrow Z$ encoding sequences into a latent space, and a decoder $D:Z\rightarrow X$ that reconstructs the sequence corresponding to a latent point. Here $Z$ is typically the latent space of a generative model. The conventional choice is to use the latent space of a VAE, where the encoded mean $\vec{\mu}_{\theta}(\vec{x})$ of a trained VAE with fitted model weights $\theta$ is used in the optimization 26. More formally, Latent Bayesian Optimization is the process of solving
$\mathop{\text{argmax}}_{\vec{x}\in X}f(\vec{x})$
@@
$\mathop{\text{argmax}}_{\vec{z}\in Z}f(D(\vec{z}))$
i.e. Bayesian Optimization over the latent space of a generative model or other invertible mapping. The surrogate model $\hat{f}:Z\rightarrow\mathbb{R}$ approximates the relationship between latent points and their corresponding objective value; for a peptide sequence $\vec{x}$ and its encoded representation $\vec{\mu}_{\theta}(\vec{x})$, the surrogate model approximates $\vec{\mu}_{\theta}(\vec{x})\rightarrow f(\vec{x})$. At iteration $l$, the acquisition function determines the subsequent point in the latent space $\vec{\mu}_{l+1}$ to decode into a peptide sequence $\vec{x}_{l+1}$ we can evaluate with the objective function $f(\vec{x}_{l+1})$. In other words, rather than performing the Bayesian Optimization on what we want to design (peptide sequences) we perform it over a (learned) representation of those sequences.
@@
In this work we are interested in finding peptide sequences with strong antimicrobial activity. We choose to use minimum inhibitory concentration (MIC) as a measure for antimicrobial activity. The MIC refers to the smallest concentration of peptide that inhibits the growth of bacteria in a standardized assay35. We make this choice because it is a theoretically continuous, standard, and meaningful measure of antimicrobial activity, and standardized data is available for measurements on it 36. Because a lower MIC corresponds to stronger antimicrobial activity, we maximize the negative of MIC: $M=-\log_{10}(MIC)$, thereby minimizing MIC.
In this pilot study, we wished to assess the impact of algorithmic choices and ensure our understanding of the design space prior to engaging in time-consuming experimental work. Therefore, we use a proxy model (also called an “oracle”) based on traditional machine learning techniques (see next section) as our “ground truth” to estimate the MIC values of sequences for which no experimentally-measured MIC is available. This is analogous to proof-of-concept small-molecule discovery projects that optimize for computed estimates of drug-likeness and solubility; a discipline-relevant proxy model is used to motivate an algorithmic exploration. In future work, we plan to leverage the procedure here to inform peptide design in more realistic contexts.
@@ 2.4 Variational Autoencoder
For the generative model whose associated latent space becomes our continuous design space of peptide sequences, we employ a variational autoencoder (VAE) based neural network. A VAE is a probabilistic deep learning architecture comprising two neural networks: the encoder $E_{\phi}:X\rightarrow Z$, which takes an input sequence and projects it to a continuous-valued latent space $z=E_{\phi}(x)$; and the decoder $D_{\theta}:Z\rightarrow X$ modules, which maps a point in the latent space back to the original input sequence space $\hat{x}=D_{\theta}(z)$ with $\phi$ and $\theta$ representing learned model weights 37. Once the model has been adequately trained, we can use the latent space and the decoder to generate new peptides, sampling from a distribution that is similar to the distribution found in the training set.
Training a VAE entails minimizing the Evidence Lower Bound (ELBO), a combination of a reconstruction loss, which measures the difference between the input sequence and the corresponding generated output sequence, and the Kullback-Leibler Divergence which functions to push the probability distribution of the latent points to be close to a Gaussian prior modeled by the latent space. We define the reconstruction loss as is typically done for sequence-based tasks 38 as the cross-entropy of the probability of predicting the next token $x_{i+1}$ conditioned on a context $\vec{c}$, usually the previous tokens in the sequence:
@@
We use a VAE with a transformer architecture as described previously 39, 25, known therein as a TransVAE. The transformer architecture is largely similar to the architecture originally proposed in Vaswani et al. 40. Input sequences were tokenized into integers then fed through a learned embedding layer. During training, sequences were predicted autoregressively, with the $(n+1)$-th token predicted from a context of the previous $n$ tokens , where the previous $n$ tokens are given directly from the training set and the reparameterized latent point representation $\vec{z}$. The embedding layer and a positional encoding layer then together feed into the encoder module consisting of self-attention blocks followed by convolution blocks. Each self-attention block is a sequence of layers: layer normalization, multi-head self-attention, dropout, a residual connection, then layer normalization, feed-forward, dropout, and another residual connection. After the self-attention blocks, there is a convolution block consisting of three layers of convolution, pooling, and leaky ReLU. The convolution block outputs the mean and variance for the VAE latent space. The decoder module consists of largely the same architecture in reverse: convolution blocks followed by masked self-attention blocks. In our previous results 25 we found evidence suggesting the TransVAE architecture may limit latent space interpretability compared to other architectures, however we did observe high performance at the assigned tasks. Here, we chose to use a transformer architecture because of their effectiveness at sequence-to-sequence tasks, and because traversing their latent space can correlate well with a predicted property 25, an important aspect of this work. Additionally, to limit the scope of this work, we focused on only the TransVAE architecture, but in future work will expand to other architectures and representation learning choices.
As previous work has shown, jointly training a property predictor can induce latent space organization 26, 25. In this work, we investigate the impact of the choice of predictive task and the availability of data associated with that task on the performance of the LBO algorithm. To do so, we train multiple TransVAE models jointly with a property predictor consisting of two feed-forward layers (see next section). The property predictor loss term $\mathcal{L}_{\text{PropPred}}$ is a simple mean-squared error between the property predictor’s outputs $y_{\text{pred}}$, and the computed property values $y_{\text{true}}$,
@@ 2.5 TransVAE Models and Training
To investigate the importance of additional information on the LBO algorithm and on the latent space that it traverses, we trained TransVAE models in conjunction with property predictors of different combinations of the physico-chemical characteristics Boman Index, hydrophobicity, and charge. Jointly training property predictors requires target values of these characteristics for corresponding (latent representations of) sequences. We computed Boman Index, net charge at a pH of 7.2, and hydrophobicity of every sequence in both the training set and test set. Boman Index is defined by summing the solubility values of individual amino acids in the peptide, normalized by length 41. It functions as an estimate of whether a peptide will bind to membranes or other proteins. For hydrophobicity, we use the Kyte-Doolittle scale 42 and average the hydrophobicity values of amino acids in a given peptide for an overall estimate of peptide affinity with water. Charge at a pH of 7.2 is simply computed by summing the predicted charge of each amino acid in the peptide at a pH of 7.2. These physicochemical properties are of interest in AMP design because of they relate to whether an AMP will interact with or penetrate a cell membrane.
@@
Overall, we studied the following models: (i) TransVAE trained jointly with a regressor directly predicting oracle values, i.e. $\log_{10}MIC$ values estimated by a predictor model; (ii) TransVAE trained jointly with a regressor predicting Boman index values; (iii) TransVAE trained jointly with a regressor predicting charge values; (iv) TransVAE trained jointly with a regressor predicting hydrophobicity values; (v) TransVAE trained jointly with a regressor predicting both Boman index and charge; and (vi), TransVAE trained jointly with a regressor predicting Boman Index, charge, and hydrophobicity values.
@@ 2.6 Dataset
To train our TransVAE models we used a peptide sequence dataset consisting of 665,772 sequences. We sourced sequences from from AMP-specific databases DRAMP V3.0 44, APD3 45, GRAMPA 36, dbAMP 2.0 46, StarPep 47, and general protein databases SwissProt and TREMBL 48. Using peptide sequences from databases of known antimicrobial peptides ensures that the data distribution — and therefore the distribution being learned by the TransVAE — contains examples of sequences that are AMPs. Additionally, to ensure the training distribution also contained non-AMPs, we excluded those peptides from SwissProt annotated with the keywords antifungal, antimicrobial, antibiotic, antiviral defense, antiviral protein, secreted; although the impact of keyword combinations on predictive model performance has been discussed elsewhere49. Lastly, we clustered TREMBL sequences at 90% identity using CD-HIT 50, 51, then randomly sampled 7500 peptides at each length, yielding a subset of 668,343 peptides (the smallest length in the TREMBL set after clustering was 15). Once we removed duplicates, the final dataset of unique peptides from every database was 665,772 sequences. This final amount was split into a training set of 523,848 sequences and a test set of 141,924 sequences where we ensured that the training set and test set had the same proportion of verified AMP and verified non-AMP peptides (SI Fig. S2).
@@ 2.8 Latent Bayesian Optimization Procedure
To perform latent BayesOpt, we use the BoTorch Python package 32, leveraging its SingleTaskGP Gaussian Process object with input unit hypercube and output standard normalization, as suggested by the codebase. As noted above, we use the Log Expected Improvement (LogEI) acquisition function, which offers improved numerical stability compared to conventional Expected Improvement 34. For each BayesOpt experiment, we perform five different optimization runs of 500 iterations (’oracle’ calls). Each run is initialized with 100 points randomly sampled from predicted-$\log_{10}$(MIC) values (see Supplementary Information for additional details). Every BayesOpt experiment uses the same 5x100 initialization points, unless otherwise specified.
@@ 3.1 Property prediction leads to latent space organization, even in data-sparse conditions
To assess the latent spaces associated with the different models, we quantified the extent to which the top five principal components of a PCA projection correlated with different physicochemical properties used in training. We then employed these PCs to visualize relevant two-dimensional slices of the latent space. To do this we first computed each property over the training set of sequences. Next, we embedded those sequences in the full latent space of each model. Then we decomposed the space into its top five principal components. For each principal component, we computed the Pearson correlation coefficient ($r$) between it and the property values.
In Fig. 2, we visualize the PC components most highly correlated with Boman Index, Charge, and Hydrophobicity for the models trained with all three. In line with previous results 26, 25, we observe latent space organization when jointly-training a decoder and a property predictor. Additionally, training a property predictor predicting multiple different properties results in an organized latent space where different properties correlate most with different pairs of dimensions (Fig. 2 rows; SI Table S5, particularly Boman index and charge). Furthermore, varying percentage of property labels — thereby limiting the amount of labels the model has access to in training — still results in latent space organization, as indicated by the persistence of an orderly color gradient between the first and last rows of Fig. 2. Overall, increasing the number of properties and masking even a large proportion of the information for the predictor did not prevent the latent space from being organized. (For a more detailed discussion of correlations between PC components and organizing variables, we refer the interested reader to Sect. S3.4 in the SI.)
Given that for 64-dimensional latent spaces we are visualizing just two-dimensional slices through PCA projections, it is possible that the organization is being induced by the projection. To check the extent to which PCA is distorting the original, high-dimensional manifold, and thus the extent to which it may be providing an illusion of organization when the high-dimensional cloud of points is not actually organized, we compute four manifold distortion metrics as we have done previously 25: trustworthiness and continuity 53, steadiness and cohesiveness 54.
@@
We compute the extent of manifold distortion between the full latent space and the 5-dimensional PCA projections using a held out test set of sequences. We observe similar levels of distortion between high-dimensional latent spaces and their PCA projections for models trained with different organizing properties and different percentages of labels (Fig. 3). Trustworthiness, continuity, and steadiness are all relatively high, with each organizing method exhibiting values >0.75>0.75 (Fig 3abc), suggesting high-quality projections with few artificial clusters (trustworthiness), minimal loss of clusters (continuity), and little stretching of distances between points (steadiness). Cohesiveness, lying near $0.6$ for all, is moderate, suggesting that compression is occurring (Fig 3d), as it necessarily must when moving from high to low dimensions. The similarity of the results and the comparatively low-to-moderate distortions observed provide confidence that the visualizations reflect extant trends between, and patterns in, the high-dimensional latent spaces.
@@
Given the number of different physicochemical properties by which a latent space may be organized, we perform an ablation study testing different numbers and combinations. We hypothesized that organizing the latent space with more properties would lead to peptide representations more amenable to BayesOpt. We summarize the performance at different iterations in Figs S10 and Figs 5a, 5b, where we investigate average best scores found under different conditions. We primarily focus on PCA5 models. We choose the top five components to balance capturing a larger amount of explained variance (near 20%, SI Fig. S4) with losing interpretability, and because three of the four scenarios where we varied the number of PCA components demonstrated five components performing better than two components (previous section).
@@
We consider single property organization first. There is no clear advantage of any single property when searching through the 64-dimensional latent spaces–the best-performing single-property models compared to other models with only organizing property changed are id-c-100%, id-b-75%, id-h-50%, id-h-2% at 100 iterations (Fig. 5a), and id-b-100% and id-c-100%, id-b-75%, id-d-50%, id-c-25%, and id-b/c-2% at 500 iterations (Fig. 5b). For PCA-projected spaces, we observe that charge-organized spaces tend to demonstrate better performance: at 100 iterations, PCA5-c-100% clearly outperforms b/h-100%, PCA5-c-75% clearly outperforms b/h-75%, PCA5-c-50% clearly outperforms b/h-50%, and PCA5-c-2% clearly outperforms b/h-2%. Only for the PCA5-25% models is charge not the clear winner, and in that case, it performs nearly as well as PCA5-h-25%. At 500 iterations, PCA5-c-100% clearly outperforms b/h-100%, PCA5-c-75% clearly outperforms b/h-75%, and PCA5-c-50% clearly outperforms b/h-50%. However, PCA-5-c-25% performs worse than b/h-25%, and all PCA5-2% models perform roughly the same, displaying rather poor performance at <0.3 average best scores. Thus, charge is a better organizing property than Boman index or hydrophobicity, particularly when more labels are available. This trend matches with a similar trend in relevancy: charge correlates more strongly and has higher mutual information with the oracle values when looking at the full peptide training set (SI Section S3.4).
@@
Considering multi-property organization, we still do not see clear trends in when searching through the full 64-dimensional latent space. After 100 iterations (Fig. 5a), id-bch-100% is one of the best performers (with performance about equivalent to id-c-100%), and id-bch-75% outperforms other models, but the highest-performing model for lower label percentages is one or more of the single-property models. After 500 iterations (Fig. 5b), we see similar model performance across the board, though id-bch-100% is one of the highest-performing models (tied with id-b-100%) and id-bc-75% is one of the highest-performing models (tied with id-b-75%) but for lower label percentages single-property models are the best, and the performance is not too dissimilar within error bars.
@@
When searching through a PCA projection, there is not a clear advantage to using multi-property organization. After 100 iterations (Fig. 5a), at 100% and 75% label percentage, charge clearly outperforms multi-property organization. At 50% label percentage, PCA5-bc-50% outperforms PCA5-c-50%, at 25% label percentage, PCA-bch-25% by far outperforms any other models, and at 2% the highest performing model is PCA5-bch-2%. After 500 iterations (Fig. 5b), at 100% and 75% label percentage, charge still outperforms multi-property organization. For lower label percentage, multi-property organization does outperform single-property organization (PCA5-bc-50% > PCA5-bch-50% > PCA5-c-50%, PCA5-bch-25% is the clear best performer, PCA5-bch-2% outperforms PCA5-bc-2% and both clearly outperform the single-property organizers), but neither bc nor bch is consistently better than the other. This suggests that multi-property organization may be better at lower label percentage, an observation which is supported by the fact that the best-performing model we observed in the study was PCA-20-oracle-2%, with a final performance of $\langle\mathcal{M_{\textrm{final}}}\rangle=$0.896 $\pm$ 0.092, where the oracle is itself a multi-property organizer, although it was not part of the ablation study.